. Let the probability distribution function of X be given by f(x) = 0.25 exp(−0.25x), x > 0 Show that E(X) = 4 and V ar(X) = 16


1
Expert's answer
2021-05-07T09:45:33-0400
"E(X)=\displaystyle\int_{0}^{\infin}x(0.25e^{-0.25x})dx"


"\int0.25xe^{-0.25x}dx"


"\int udv=uv-\int vdu"

"u=0.25x, du=0.25dx"


"v=-4e^{-0.25x}"



"\int0.25xe^{-0.25x}dx=-xe^{-0.25x}+\int e^{-0.25x}dx"

"=-xe^{-0.25x}-4e^{-0.25x}+C"

"E(X)=\lim\limits_{t\to \infin}\displaystyle\int_{0}^{t}x(0.25e^{-0.25x})dx"

"=\lim\limits_{t\to \infin}[-xe^{-0.25x}-4e^{-0.25x}]\begin{matrix}
 t \\
 0
\end{matrix}=4"



"\int0.25x^2e^{-0.25x}dx"


"\int udv=uv-\int vdu"

"u=0.25x^2, du=0.5xdx"


"v=-4e^{-0.25x}"



"\int0.25x^2e^{-0.25x}dx=-x^2e^{-0.25x}+2\int xe^{-0.25x}dx"

"=-x^2e^{-0.25x}-8xe^{-0.25x}-32e^{-0.25x}+C"


"E(X^2)=\lim\limits_{t\to \infin}\displaystyle\int_{0}^{t}x^2(0.25e^{-0.25x})dx"

"=\lim\limits_{t\to \infin}[-x^2e^{-0.25x}-8xe^{-0.25x}-32e^{-0.25x}]\begin{matrix}
 t \\
 0
\end{matrix}"

"=32"


"Var(X)=E(X^2)-(E(X))^2"


"=32-(4)^2=16"